12. A mail order company classifies its customers by gender and location of residence

Table for Exercise 12 - Mail order Customers

Location

Gender

Male

Female

Suburban

196

298

Urban

92

173


a. What is the probability that a customer is male?

b. What is the probability that a customer is female?

c. What is the probability that a customer is a suburban male?

d. What is the probability that a customer is an urban female?

e. What is the probability that a customer is a suburban male or an urban female?

f. What is the probability that a customer is urban?

g. What is the probability that a customer is not a suburban female?

h. What approach to probability did you use to calculate your answers?

i. Are the events {customer is urban} and {customer is suburban} mutually exclusive?

Explain.
Answer #1

Answer)

Probability is given by favorable/total

Total = 196 + 298 + 92 + 173 = 759

A)

Favorable = male = 196 + 92 = 288

Required probability is 288/759 = 0.37944664031

B)

Favorable = (298+173)

Required probability is (298+173)/759 = 0.62055335968

C)

Favorable = suburban male = 196

Required probability is 196/759 = 0.25823451910

D)

Favorable = urban female = 173

Required probability is 173/759 = 0.22793148880

E)

Suburban male or urban female = 196 + 173

Required probability is (196+173)/759 = 0.48616600790